# suppSupplementary References

### v1

@@ Abstract
We propose a method, HotSpot, for optimizing neural signed distance functions, based on a relation between the solution of a screened Poisson equation and the distance function. Existing losses such as the eikonal loss cannot guarantee the recovered implicit function to be a distance function, even when the implicit function satisfies the eikonal equation almost everywhere. Furthermore, the eikonal loss suffers from stability issues in optimization and the remedies that introduce area or divergence minimization can lead to oversmoothing. We address these challenges by designing a loss function that when minimized can converge to the true distance function, is stable, and naturally penalize large surface area. We provide theoretical analysis and experiments on both challenging 2D and 3D datasets and show that our method provide better surface reconstruction and more accurate distance approximation.
Figure 1: We propose HotSpot, a neural signed distance function optimization method that can obtain accurate signed distance function for complex shapes. Here we show a reconstruction from a point cloud sampled from the reference bunny (taken from Mehta et al. [1]) on the right. In the inset, we visualize the recovered signed distance function on a horizontal slice, using warm colors for positive values and cold for negative (zoom in for details). Our reconstruction is significantly more accurate than prior works (SAL [2], DiGS [3], and StEik [4]).
@@ 1 Introduction
Learning neural signed distance function requires designing loss functions so that the zero crossings of the implicit function is at the desired locations (e.g., adhere to a sparse input point cloud), and that the implicit function returns correct signed distance. The design of the loss proves to be challenging for shapes with complex topology and details (Fig. 1). We propose HotSpot, a method based on a relation between screened Poisson equation and distance that enjoys both theoretical and practical benefits.
@@
A major challenge of signed distance function optimization is to ensure that the implicit function indeed outputs the signed distance. A standard regularization loss used is the eikonal equation: it constrains the norm of the gradient of an implicit function to be $1$ almost everywhere. If the implicit function is a signed distance function, then it satisfies the eikonal equation. However, the converse is not true. Fig. 2 shows an example: on the left, we optimize an implicit function to satisfy the eikonal equation, while it successfully does so, it converges to a solution that is far from the actual distance [5, 6].
Figure 2: We illustrate an 1D neural signed distance function optimization using the classical eikonal loss [7] and our model. The $x$ axis is the domain and the $y$ axis shows the output of the impilcit function. The gray dashed line is the reference signed distance function. The middle and bottom rows show the intermediate and final states of the optimization respectively. The eikonal loss is incapable of converging to the actual signed distance function (dashed line), even when the function satisfies the eikonal equation almost everywhere. Our method converges to the ground truth by modeling a screened Poisson equation.
@@
Another key challenge lies in the optimization. It is known that the eikonal equation regularization is unstable [4], and the instability can make optimization converge to suboptimal results or even diverge. Furthermore, to combat both the ill-posedness and the instability of the optimization, existing methods need to penalize the surface area of the shape, which leads to a delicate balance between over- and under-smoothing.
@@
We propose a simple model that can be used together with the eikonal equation to address all the challenges above. Our model, HotSpot, is based on a screened Poisson equation and a classical relation between heat transfer and distance [8], that was popular for approximating both Euclidean and geodesic distance [9, 10]. Using this relation, we design a loss function for optimizing neural signed distance functions. As our contributions, we show, both theoretically and empirically, that our model
  • has a bounded difference from the actual distance and can converge to it (Fig. 2, right).
@@
  • is stable both spatially and temporally, in the sense that a slight perturbation of the implicit function only introduce changes in a local region, and that the dynamical system formed by the gradient flow will converge in the long run.
@@
  • penalizes the surface area naturally, while being compatible with being a signed distance function.
@@
  • works well in both 2D and 3D shapes, and it excels at approaching the signed distance functions while optimizing for complex, high genus shapes.
@@ 2 Related Work
Implicit surfaces. Modeling and reconstructing surfaces using implicit functions has a long history in computer vision and graphics [11, 12, 13, 14, 15]. Recently, there is a surge of interests to model surfaces using neural implicit functions [16, 17, 18, 19, 20]. Neural implicit representations are favored for their compactness, ability to represent highly-detailed surfaces, and compatibility to gradient-based optimization [1].
@@
Neural signed distance functions. Among the classes of implicit surfaces, signed distance functions, where the implicit function outputs the distance to the closest point on the surface, and the sign represents whether the point is inside or outside the solid object, is one of the most popular variants. Signed distance functions enable an efficient rendering algorithm called sphere tracing [21, 22, 23, 24] and helps geometry compression [25]. They can also be used for collision detection [26, 27, 28], and numerical simulation [29]. Traditionally, signed distance functions are often represented as grids [6]. More recently, neural signed distance functions have become an appealing geometric representation [30, 31, 32]. (Neural) Signed distance functions are also often used in inverse rendering [33, 34, 35, 36, 37, 38, 39], for their flexibility to obtain surfaces with complex topology.
@@
Regularizing and initializing neural signed distance functions. A commonly used loss for optimizing neural signed distance functions is the eikonal regularization, which encourages the norm of the gradient of the implicit function to be one [7, 36]. However, the eikonal regularization alone is far from sufficient, as 1) it can contain many undesirable local minima [4], and 2) minimizing eikonal loss does not necessarily lead to a distance field [5, 6]. Previous works have proposed initialization to bootstrap the optimization with a simple shape [7, 3]. Recent works show that minimizing the divergence and the surface area of the implicit function can lead to better reconstruction [3, 40, 41] and stabilizes the optimization [4]. Second-order information is also often used [41, 42] for regularization. Specialized architectures have been proposed for limiting the Lipchitz constant of the implicit function [32] and for representing high-frequency details [31, 25, 43]. Lipman [44] showed a connection between the regularization of occupancy function optimization and an approximated signed distance field using the phase transition theory of fluids.
@@
We propose a new loss and method for optimizing signed distance functions. While derived from very different mathematics, our loss turns out to have a close connection to Lipman’s PHASE loss. We show that our theory leads to significantly more numerically stable optimization, and a different conclusion in the parameter setting that leads to drastically different results.
@@
Distance approximation using heat transfer. Our derivation is inspired by a relation between heat transfer and approximated distance [8]. This relation has been used in computer graphics for efficiently computing the geodesic distance on surfaces [9, 10, 45, 46]. We apply the relation for optimizing neural signed distance functions, and analyze the properties of the resulting optimization in this context.
@@ 3 Background and Motivation
However, as mentioned earlier, simply optimizing for the two terms above can lead to unsatisfactory results, as the eikonal loss does not guarantee that $u$ would be a distance function even when satisfied almost everywhere (Fig. 2), and the extra regularization previous work added to minimize the area and divergence [3, 4] could further tamper with the distance approximation quality. We derive an additional loss that can be used together with both of the losses above, and analyze our loss theoretically and empirically.
(a) 1D Screened Poisson Equation
@@
Figure 3: We show an illustration of the relation between screened Poisson equation and a distance field. The top shows an 1D example, where on the left we show a solution to Eq. 4 with different absorption $\lambda$, and on the right we show the reconstructed distance $\frac{1}{\lambda}\ln(h)$ (Eq. 5). As $\lambda$ increases, the heat cools down faster, and the reconstructed function approaches the real distance. The bottom shows a 2D example.
@@
Fig. 3 shows the intuition. Screened Poisson equation simulates heat transfer until an equilibrium is reached, under absorption controlled by the $\lambda$ coefficient. As we increase the absorption, the reconstructed distance $\frac{1}{\lambda}\ln(h_{\lambda})$ would approach the real distance. In practice, we could only obtain a discrete point set $\Gamma_{0}$. Provided that the neighbor points are close enough, if there are some boundaries connecting them, the boundary condition of Eq. 4 still holds. This can be interpreted as a superset $\Gamma$ replacing $\Gamma_{0}$, effectively transforming the boundary from a discrete point cloud into a continuous surface, which is required and precisely the goal in the signed distance function reconstruction. The reconstructed distance can still faithfully represent the distance to $\Gamma$ because the approximation in Eq. 5 still holds. We will later prove that the measure of $\Gamma$ is bounded by our loss.
@@ 4 Method
In this section, we introduce our method, HotSpot, based on the screened Poisson equation (Eq. 4) and derive theoretical analysis. We use the screened Poisson equation to derive a loss we can directly apply to an implicit function $u$, and prove that there is a bound to the solution of the screened Poisson equation to the true distance (Section 4.1). We further derive stability analysis of our loss on both the spatial stability over small perturbation and the temporal convergence (Section 4.2). Next, we show that our loss can penalize surfaces with large area (Section 4.3). Finally, we discuss the relation to a prior work, PHASE (Section 4.4).
@@ 4.1 Screened Poisson Equation Informed Signed Distance Functions Optimization
We want to design a loss such that its minimizer would satisfy the screened Poisson equation. This can be achieved by minimizing the energy functional $\frac{1}{2}\int_{\Omega}\|\nabla h(\bm{x})\|^{2}+\lambda^{2}h(\bm{x})^{2}\,\mathrm{d}\bm{x}$. After substitution and removal of the constant factor $\lambda$, we obtain our heat loss:
@@
If we take the derivative of $L_{\text{heat}}$ with respect to the heat field $h$, we can see that it recovers the screened Poisson equation $\nabla^{2}h-\lambda^{2}h=0$. The boundary condition can be enforced by the standard boundary loss (Eq. 2), by observing that $u(\bm{x})=0$ implies $h(\bm{x})=1$. Together with the eikonal loss, we optimize:
@@
While prior work [8] has studied the limiting behavior of the absorption coefficient $\lambda$, to our knowledge, the bounds of the approximation and the convergence speed remains unknown. In the supplementary material A.3, we prove that, for a boundary $\Gamma$ formed by a discrete set of points and a small volume around them, the distance approximation obtained from the solution of the screened Poisson equation converges linearly with respect to $\lambda$:
@@
Here, both $C_{\text{upper}}(\bm{x})$ and $C_{\text{lower}}(\bm{x})$ are scalar fields independent to $\lambda$. The inequalities are tight, which means that one can find a case where the equality holds.
Figure 4: We visualize the direction of the changes of the implicit function during one iteration of optimization. Left shows the result with only the eikonal loss and right shows the result after adding our loss. Our loss avoids the local minimum in Fig. 2.
Compared to the eikonal loss (Eq. 3), where the error to the actual distance can be unbounded even when the implicit function satisfies the eikonal equation almost everywhere, Eq. 9 highlights the fact that even with a finite absorption $\lambda$, our loss can lead to much more accurate distance. In fact, any loss that only depends on the norm of the gradients $\left\|\nabla u\right\|$ would be incapable of distinguishing between a pseudo signed distance function [5] and an actual one, due to the discontinuous jumps of the gradient. Adding our loss (Eq. 7) ensures that the implicit function converges to the actual signed distance. Fig. 4 visualizes the direction of how an implicit function would evolve for only using the eikonal loss and after adding our loss. Our loss ensures that the implicit function evolves towards the direction of the actual distance.
@@ 4.2.1 Spatial Stability
We show that for the screened Poisson equation, an error introduces to a single point in the solution decays exponentially and do not propagate to infinitely distant positions. On the contrary, for the corresponding partial differential equation formed by the eikonal equation, the error can propagate to infinitely far away. The property of the screened Poisson equation ensures that the perturbation we introduce during optimization does not lead to a drastically different target and stablizes the optimization.
First, we show that for the partial differential equation formed by the eikonal equation, an additive error introduced to a point can preserve along a line segment or a ray.
@@
The proofs of these two propositions are in the supplementary material A.1 and A.2. By comparing the error fields, we observe that the partial differentiation equation governed by the screened Poisson equation exhibits greater spatial robustness than the one based on eikonal equation.
@@ 4.2.2 Temporal Stability
We next demonstrate the temporal stability of the heat loss and show that the dynamical system of the gradient update is stable, that is, the solution of the dynamical system would converge as the time $t$ approaches infinity. This analysis is supported by von Neumann stability analysis, similar to Yang et al.’s work [4].
@@
They analyze the stability by applying Fourier transform to both sides:
@@
This analysis reveals that the original process $h(\bm{x},t)$ converges over time and remains stable. Due to the constraint $|u|=-\ln(h)/\lambda$, the term $|u(\bm{x},t)|$ also converges during the optimization process. Since $u$ is continuous, it converges as well.
Based on this analysis, we have demonstrated that our method exhibits greater stability compared to the eikonal equation in both the spatial and temporal domains.

### v2

@@ Abstract
We propose a method, HotSpot, for optimizing neural signed distance functions. Existing losses, such as the eikonal loss, act as necessary but insufficient constraints and cannot guarantee that the recovered implicit function represents a true distance function, even if the output minimizes these losses almost everywhere. Furthermore, the eikonal loss suffers from stability issues in optimization. Finally, in conventional methods, regularization losses that penalize surface area distort the reconstructed signed distance function. We address these challenges by designing a loss function using the solution of a screened Poisson equation. Our loss, when minimized, provides an asymptotically sufficient condition to ensure the output converges to a true distance function. Our loss also leads to stable optimization and naturally penalizes large surface areas. We present theoretical analysis and experiments on both challenging 2D and 3D datasets and show that our method provides better surface reconstruction and a more accurate distance approximation.
Figure 1: We propose HotSpot, a neural signed distance function optimization method that establishes an asymptotic sufficient condition to guarantee convergence to a true distance function, enabling precise surface reconstruction and level set representation for complex shapes. Here we show a reconstruction from a point cloud sampled from the reference bunny (taken from Mehta et al. [1]) on the right. In the inset, we visualize the recovered signed distance function on a horizontal slice, using warm colors for positive values and cool for negative (zoom in for details). Our reconstruction is significantly more accurate than prior works (SAL [2], DiGS [3], and StEik [4]).
@@ 1 Introduction
Learning neural signed distance function requires designing loss functions so that the zero crossings of the implicit function are at the desired locations (e.g., adhere to a sparse input point cloud), and that the implicit function returns the correct signed distance. Designing the loss function proves to be challenging for shapes with complex topology and details (Fig. 1). We propose HotSpot, a method based on a relation between the screened Poisson equation and distance that enjoys both theoretical and practical benefits.
@@
A major challenge in signed distance function optimization is ensuring the implicit function follows the true distance. A common regularization loss used is the eikonal equation, which constrains the gradient norm of an implicit function to be $1$ almost everywhere. If the implicit function is a signed distance function, then it satisfies the eikonal equation. However, the converse is not true. That is, satisfying the eikonal equation is only a necessary condition for the signed distance function, not a sufficient condition [5]. Fig. 2 presents the logic and Fig. 3 shows an example.
@@
Another key challenge lies in the optimization. It is known that the eikonal equation regularization is unstable [4], and the instability can make optimization converge to suboptimal results or even diverge.
Furthermore, to address optimization ill-posedness and instability, existing methods penalize the surface area of the shape. We prove this degrades the reconstruction quality and distorts the output level sets, thereby leading to a delicate balance between over- and under-smoothing.
@@
We propose a simple model that can address all the challenges above. Our model, HotSpot, is based on a screened Poisson equation and a classical relation between heat transfer and distance [6] that was popular for approximating both Euclidean and geodesic distance [7, 8, 9]. Using this relation, we design a loss function for optimizing neural signed distance functions. As our contributions, we show, both theoretically and empirically, that our model
  • serves as a sufficient condition ensuring the difference between the minimizer and the actual distance is bounded and decreases linearly as the parameter $\lambda$ increases (Fig. 2 and Fig. 3, right).
@@
  • is stable both spatially and temporally, in the sense that a slight perturbation of the implicit function only introduces changes in a local region, and that the dynamical system formed by the gradient flow will converge in the long run.
@@
  • penalizes the surface area naturally, while remaining an undistorted signed distance function.
@@
  • works well in both 2D and 3D shapes, and it excels at approaching signed distance functions while optimizing for complex, high-genus shapes.
Figure 2: Existing constraints are only necessary but insufficient conditions. That is, there exist solutions, like Fig. 3, that satisfy them almost everywhere yet fail to be distance functions. In contrast, we use a heat loss, modeled by a screened Poisson equation with parameter $\lambda$. As the parameter $\lambda$ goes to infinity, the minimizer converges to the distance function and excludes other solutions. Thus, our method is asymptotically both necessary and sufficient.
Figure 3: We illustrate a 1D neural signed distance function optimization using the classical eikonal loss [10] and our model. The $x$ axis is the domain and the $y$ axis shows the output of the implicit function. The middle and bottom rows show the intermediate and final states of the optimization respectively. The eikonal loss, as a necessary but insufficient condition, is incapable of converging to the actual signed distance function (dashed line), even when the function satisfies the eikonal equation almost everywhere.
@@ 2 Related Work
Implicit surfaces. Modeling and reconstructing surfaces using implicit functions has a long history in computer vision and graphics [11, 12, 13, 14, 15]. Recently, there is a surge of interest to model surfaces using neural implicit functions [16, 17, 18, 19, 20]. Neural implicit representations are favored for their compactness, ability to represent detailed surfaces, and compatibility to gradient-based optimization [1].
@@
Neural signed distance functions. Among the classes of implicit surfaces, signed distance functions, where the implicit function outputs the distance to the closest point on the surface, and the sign represents whether the point is inside or outside the solid object, is one of the most popular variants. Signed distance functions enable an efficient rendering algorithm called sphere tracing [21, 22, 23, 24] and helps geometry compression [25]. They can also be used for collision detection [26, 27, 28] and numerical simulation [29]. Traditionally, signed distance functions are often represented as grids [30, 31, 32]. More recently, neural signed distance functions have become an appealing geometric representation [33, 34, 35]. (Neural) Signed distance functions are also often used in inverse rendering [36, 37, 38, 39, 40, 41, 42], for their flexibility to obtain surfaces with complex topology.
@@
Regularizing and initializing neural signed distance functions. A commonly used loss for optimizing neural signed distance functions is the eikonal regularization, which encourages the norm of the gradient of the implicit function to be one [10, 39]. However, the eikonal regularization alone is far from sufficient, as 1) minimizing eikonal loss does not necessarily lead to a distance field [5, 30], and 2) its optimization is unstable [4]. Previous works have proposed initialization to bootstrap the optimization with a simple shape [10, 3]. Recent works show that minimizing the divergence and the surface area of the implicit function can lead to better reconstruction [3, 43] and stabilizes the optimization [4]. Second-order information is also often used [43, 44, 45, 46] for regularization. All losses above are based on necessary properties of real distance but neither establish convergence nor exclude non-SDF solutions.
Specialized architectures have been proposed to limit the Lipschitz constant of the implicit function [35], satisfy $p$-Poisson equation [47], or represent high-frequency details [34, 25, 48]. Some research [49, 50] has shown a link between occupancy function optimization and an approximate signed distance field regularization. Our experiments further reveal that some implementations [2, 51, 52] optimizing SDFs with the closest point may struggle when querying near the interpolated surface.
@@
We propose a new loss and method for optimizing signed distance functions. While derived from very different mathematics, our loss turns out to have a close connection to Lipman’s PHASE loss [49]. We show that our theory leads to significantly more numerically stable optimization, and a different conclusion in the parameter setting that leads to drastically different results.
@@
Distance approximation using heat transfer. Our derivation is inspired by a relation between heat transfer and approximated distance [6]. This relation has been used in computer graphics for efficiently computing the geodesic distance on surfaces [7, 8, 9, 53, 54]. We apply the relation for optimizing neural signed distance functions, and analyze the properties of the resulting optimization in this context.
@@ 3 Background and Motivation
However, as noted, simply optimizing for the two terms above does not guarantee that $u$ would be a distance function even when satisfied almost everywhere (Figs. 2 and 3), and the extra regularization previous work added to minimize the area and divergence [3, 4] could further tamper with the distance approximation quality. We derive an additional loss that can be used together with both of the losses above, and analyze our loss theoretically and empirically.
(a) 1D Heat Simulation via Eq. 4
@@
Figure 4: We show an illustration of the relation between screened Poisson equation and a distance field. The top shows a 1D example, where on the left we show a solution to Eq. 4 with different absorption $\lambda$, and on the right we show the reconstructed distance $-\frac{1}{\lambda}\ln(h)$ (Eq. 5). Boundary loss for real boundary points acts as an isothermal heat source, while heat loss diffuses the heat. When $\lambda$ is increased, the heat decays faster, and the error of the reconstructed function approaches 0. The bottom shows a 2D example.
@@
Fig. 4 shows the intuition. The screened Poisson equation simulates heat transfer until an equilibrium is reached, under absorption controlled by the $\lambda$ coefficient. As we increase the absorption, the reconstructed distance $\frac{1}{\lambda}\ln(h_{\lambda})$ would approach the real distance. In practice, we could only obtain a discrete point set $\Gamma_{0}$. Provided that the neighbor points are close enough, if there are some boundaries connecting them, the boundary condition of Eq. 4 still holds. This can be interpreted as a superset $\Gamma$ replacing $\Gamma_{0}$, effectively transforming the boundary from a discrete point cloud into a continuous surface, which is required and precisely the goal of the signed distance function reconstruction. Obtaining a distance function to the point cloud rather than the interpolated surface is not our objective. The reconstructed distance can still faithfully represent the distance to $\Gamma$ because the approximation in Eq. 5 still holds. Later, we will prove that the measure of $\Gamma$ is bounded by our loss.
@@ 4 Method
In this section, we introduce our method, HotSpot, based on the screened Poisson equation (Eq. 4) and derive a theoretical analysis. We use the screened Poisson equation to derive a loss we can directly apply to an implicit function $u$, and prove that there is a bound to the solution of the screened Poisson equation to the true distance (Section 4.1). We further derive a stability analysis of our loss on both the spatial stability over small perturbations and the temporal convergence (Section 4.2). Next, we show that our loss can penalize surfaces with large area (Section 4.3). Finally, we discuss the relation to prior work, PHASE (Section 4.4).
@@ 4.1 Screened Poisson Equation Informed Signed Distance Functions Optimization
We want to design a loss such that its minimizer satisfies the screened Poisson equation. This can be achieved by minimizing the energy functional $\frac{1}{2}\int_{\Omega}\|\nabla h(\bm{x})\|^{2}+\lambda^{2}h(\bm{x})^{2}\,\mathrm{d}\bm{x}$. After substitution and removing of the constant factor $\lambda$, we obtain our heat loss:
@@
If we take the derivative of $L_{\text{heat}}$ with respect to the heat field $h$, we recover the screened Poisson equation $\nabla^{2}h-\lambda^{2}h=0$. The boundary condition can be enforced by the standard boundary loss (Eq. 2), by observing that $u(\bm{x})=0$ implies $h(\bm{x})=1$. Together with the eikonal loss, we optimize:
@@
While prior work [6] has studied the limiting behavior of the absorption coefficient $\lambda$, to our knowledge, the bounds of the approximation and the convergence speed remain unknown. In the supplementary material A.3, we prove that, for a boundary $\Gamma$ formed by a discrete set of points and a small volume around them, the distance approximation obtained from the solution of the screened Poisson equation converges linearly with respect to $\lambda$:
@@
Here, both $C_{\text{upper}}(\bm{x})$ and $C_{\text{lower}}(\bm{x})$ are scalar fields independent of $\lambda$. The inequalities are tight, which means that one can find a case where the equality holds.
Figure 5: We visualize the direction of the changes of the implicit function during one iteration of optimization. Left shows the result with only the eikonal loss and right shows the result after adding our loss. Our loss helps to escape the local minimum in Fig. 3. The heat diffusion encourages the reconstructed absolute distance to increase monotonically away from the boundary.
Compared to the eikonal loss (Eq. 3), where the distance error can be unbounded even when the equation holds almost everywhere, Eq. 9 shows that even with a finite absorption $\lambda$, our loss can lead to much more accurate distance. In fact, any loss that only depends on the norm of the gradients $\left\|\nabla u\right\|$ is not a sufficient condition of the distance function (Fig. 2), due to the discontinuous jumps of the gradient. Adding our loss (Eq. 7) ensures that the absolute value of the implicit function converges to the actual distance. Fig. 5 visualizes the direction of how an implicit function would evolve when using only the eikonal loss and after adding our loss. Our loss ensures that the implicit function evolves towards the actual distance.
@@ 4.2.1 Spatial Stability
We show that for the screened Poisson equation, an error introduced to a single point in the solution decays exponentially and does not propagate to infinitely distant positions. On the contrary, for the eikonal equation, the error can propagate to infinitely far. The property of the screened Poisson equation ensures that the perturbation introduced during optimization does not lead to a drastically different target and stabilizes the optimization.
First, we show that for the eikonal equation, an additive error introduced to a point can presist along a line segment or a ray.
@@
The proofs of these two propositions are in the supplementary material A.1 and A.2. By comparing the error fields, we observe that the partial differential equation governed by the screened Poisson equation exhibits greater spatial robustness than the one based on the eikonal equation.
@@ 4.2.2 Temporal Stability
We next demonstrate the temporal stability of the heat loss and show that the dynamical system of the gradient update is stable, that is, the solution of the system converges as the time $t$ approaches infinity. This is supported by von Neumann stability analysis, similar to Yang et al.’s work [4].
@@
They analyze the stability by applying Fourier transform:
@@
This analysis reveals that the original process $h(\bm{x},t)$ converges over time and remains stable. Due to the constraint $|u|=-\ln(h)/\lambda$, the term $|u(\bm{x},t)|$ also converges during the optimization process. Since $u$ is continuous, it converges as well. Thus, our method exhibits greater spatial and temporal stability compared to the eikonal equation.